The table below represents how Marco feels about chocolate candy bars.

a. Fill in the missing values for total utility and marginal utility.

Instructions: Enter your answers as a whole number.

Chocolate Candy Bars and Marco's Utility

Chocolate Candy Bars Total Utility (utils) Marginal Utility (utils)
0 0
1 26
2 18
3 57
4
5 71 5
6 0

Suppose Marco currently has two candy bars. You tell Marco you will give him either a soda, which gives him 26 utils of happiness, or two additional candy bars.

b. Which is he likely to prefer?

         (Click to select)   A soda   Two extra candy bars   
Answer #1

Solution:-

Quantity

Total Utility

Marginal Utility

0

0

-

1

26

26

2

(26 + 18) = 44

18

3

57

(57 – 44) = 13

4

(71 – 5) = 66

(66 – 57) = 9

5

71

5

6

71

0

Since Macro has two chocolate candy bars, so total utility from this is 44. If he gets two additional chocolate candy bars, then her total utility becomes 66. Instead if he gets soda then his total utility becomes 44+26 = 70.

So he is likely to prefer a soda.